OpenStudy (lollola):

A researcher tagged 40 giant turtles in 2008. In 2009 she returned and captured 50 turtles, 16 of which were tagged. in 2010 she returned again and captured 30 turtles 8 of which were tagged. a: Use a proportion to estimate the turtle population for 2009 b: Use a proportion to estimate the turtle population for 2010 Can someone help me? I really dont understand what to do, so as simply as you can explain it the better. Ty <3

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (lollola):

Well I have never worked with these types of problems before so i'm not even sure how to set up the equation.

OpenStudy (skullpatrol):

Do you know what a proportion is?

OpenStudy (skullpatrol):

A proportion is an equation that states that two ratios are equal.

OpenStudy (lollola):

ok

OpenStudy (skullpatrol):

Generally a/b = c/d

OpenStudy (skullpatrol):

40/b = 16/50

OpenStudy (lollola):

so we are finding b?

OpenStudy (skullpatrol):

Yes.

OpenStudy (lollola):

How would we find b?

OpenStudy (phi):

you are allowed to "flip" *both* ratios, so do that: b/40= 50/16 now multiply both sides by 40

OpenStudy (lollola):

I confused. Should I solve 50/16 before multiplying?

OpenStudy (lollola):

so .. b/40= 12.8?